Suppose that the sales function for a product A has been estimated as Qu=71.6R-2Pct.8Ps+.000 5Y+·0004A Where, Qu is the quantity of product L demanded, in thousand per month. PL is the price of product L, which is currently $15 Pc is the price of product C, which is currently $25; Ps is the price of product S, which is currently $20; Y is the level of per capita income, which is currently $20,000, and A is monthly advertising, which is currently $15,000. Express the equation for the demand curve for L. Select one: a. PL-21.25-625Q b.QL#34-1.6PL 0 O d. Pc 100-5Pc
Answer #1

b. QL = 34 - 1.6PL

QL = 7 - 1.6PL - 0.2PC + 0.8PS + 0.0005Y + 0.0004A
PL = $15, PC = $25, PS = $20, Y = $20,000, A = $15,000

We substitute the values of PC, PS, Y, and A in QL.

QL = 7 - 1.6PL - 0.2(25) + 0.8(20) + 0.0005(20,000) + 0.0004(15,000)
So, QL = 7 - 1.6PL - 5 + 16 + 10 + 6
So, QL = 34 - 1.6PL